shift summary for evening 7/2/99

Theodoros Alexopoulos (theoalex@fnal.gov)
Sat, 3 Jul 1999 00:33:25 -0500 (CDT)

Shift Summary for evening shift, 7/2/99

On Shift: MT, TA
Other People present PSH, ST, RSK

Hours of beam available at all intensities: ~6.5h
Hours of beam available at nominal intensity: ~6.5h
Hours of E832 data (special runs not included): ~5.5h
Hours of special runs (laser scans, muon runs...): ~0.5h

Accounting of lost time:
starting/stopping runs,
special BEAM_832_NOL3, and BEAM_PED_NOKTEV1 runs

========================================================================
Runs ended this shift:
Run 13794 - 3.6M events - BEAM_832_NOKTEV1
Run 13796 - no tape - BEAM_832_NOL3
Run 13797 - no tape - BEAM_PED_NOKTEV1
Run 13798 - no tape - BEAM_832_NOL3
Run 13799 - still going on - BEAM_832_NOKTEV1

Tapes brought to FCC this shift:
QKS076,QKS077,QKS078,
QKC076,QKC077,QKC078,
QKK076,QKK077,QKK078,
QKB026

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Smooth E832_NOKTEV1 datataking, but beam was intermittent for
good part of the shift.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13794 16:00 run in progress, w/ BEAM_832_NOKTEV1 configuration,
since ktev1 is dead.
The SGI engineer goes and comes back with several
big boxes...
13796 17:00 start this run with DC TDC offsets shifted
by -100ns to study early hits in L3 filter. This
shift was the wrong one.
17:35 SL,VJ and AT found out that the regenerator Relay Rack
is mising the AC voltage cheaters.
Install new ones---> Ped RMSs look much
narrower now.
3 new bottles of Ar/Ethane now online westside.
17:25 Took a Ped run, 13797.
13798 17:39 start this run with DC TDC offsets shifted
by 100ns to study early hits in L3 filter. This
shift was in the right direction.
13799 17:48 Back to normal BEAM_832_NOKTEV1 configuration.
18:04 Beam goes away due to Linac quad trip.
Experts enroute.
18:25 Beam is back, but intermittent.
19:30 SGI engineer back with 3 more bigger boxes.
He is working really hard...

20:45 SGI guy fixes ktev1. He replaces OLS,
EBUS controller, and the entire back plane that
was completely damaged. ktev1 is a brand new
computer...
22:36 TeV aborts beam for several spills
23:00 Beam is back.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Keep taking data with BEAM_832_NOKTEV1 configuration overnight